FTIR OBSERVATION AND DFT STUDY OF CoC_{3} TRAPPED IN SOLID Ar

\underline{SARAH A. BATES}, J. A. RHODES, C. M. L. RITTBY, and W. R. M. GRAHAM, Department of Physics and Astronomy, Texas Christian University, Fort Worth, TX 76129.

Fourier transform infrared (FTIR) and density functional theory (DFT) isotopic studies have been initiated on cobaltcarbon species and have resulted in the detection of linear CoC_{3}. The CoC_{3} chain was produced by the dual laser ablation of carbon and cobalt rods, followed by trapping the products in solid Ar at ∼10 K. FTIR measurements of ^{13}C isotopic shifts are in good agreement with DFT predictions using both the BPW91 and B3LYP functionals with a 6-311+G(3df ) basis set, confirming the assignment of the ν_{1}(σ) fundamental of linear CoC_{3} at 1918.2 cm^{−1}.